# Groups Acting Acylindrically on Trees

### v5

@@ 2.3. Acylindrically Arboreal Groups
We say that a group $G$ is acylindrically arboreal if $G$ acts acylindrically non-elementarily on some tree $T$. Up to subdivision of edges such an action may be assumed to be without inversion, so as in [stilwell2002trees, Section I.5.4] it will give rise to a quotient graph of groups decomposition of $G$, which we will call a non-elementary acylindrical splitting of $G$.
Any acylindrically arboreal group is acylindrically hyperbolic, as any simplicial tree is a 00-hyperbolic metric space, but working with a simplicial tree allows us to use a much more combinatorial definition.
@@
[Weidmann07, Introduction] Let $G$ be a group acting on some tree $T$ and let $k\geq 0$ and $C>0$ be integers. We say that the action of $G$ on $T$ is $(k,C)$-acylindrical if the pointwise stabiliser of any edge path in $T$ of length at least $k$ contains at most $C$ elements.
The following theorem is essentially due to Minasyan and Osin [minasyanOsin13]. We include a brief proof demonstrating how to use the cited lemma to obtain this result.
@@
[minasyanOsin13, Lemma 4.2] Let $G$ be a group acting by isometries on a tree $T$. This action is acylindrical (in the sense of Definition 2.1) if and only if there exist constants $k\geq 0$ and $C\geq 1$ such that the action of $G$ on $T$ is $(k,C)$-acylindrical.
@@
For the other direction assume that the action of $G$ on $T$ is $(k,C)$-acylindrical for some constants $k$ and $C$. Let $\epsilon>0$, and set $R(\epsilon)=k+2\epsilon+6$, $N(\epsilon)=2(2\epsilon+1)C$.
@@
so it follows that
$d_{T}(u,v)\geq k\text{ and }d_{T}(\{u,v\},\{x,y\})>\epsilon,$
and we can invoke [minasyanOsin13, Lemma 4.2] to see that the pointwise quasi-stabiliser $\operatorname{PQStab}_{G}^{\epsilon}(\{x,y\})$ is contained in at most $2(2\epsilon+1)$ cosets of $\operatorname{PStab}_{G}(\{u,v\})$, the pointwise stabiliser of $\{u,v\}$. By construction the geodesic path between $u$ and $v$ must use at least $k$ edges and so $|\operatorname{PStab}_{G}(\{u,v\})|\leq C$ by definition of definition of $k$ and $C$. It follows that $|\operatorname{PQStab}_{G}^{\epsilon}(\{x,y\})|\leq 2(2\epsilon+1)C=N(\epsilon)$. Thus our action was acylindrical with constants $R(\epsilon)$ and $N(\epsilon)$. ∎
Theorem 2.17 does not guarantee that $G$ is acylindrically arboreal, only that the action of $G$ on $T$ is acylindrical. The class of acylindrically arboreal groups has the following set of inheritance properties, claim (3) of which follows immediately from a restriction of of Lemma 7.1 of [Osin13] to actions on trees.
@@
    [Osin13, Lemma 7.1] Let $H\leq G$ be s-normal, i.e. for all $g\in G$, $gHg^{-1}\cap H$ is infinite. Then $H$ is acylindrically arboreal. In particular, any finite index subgroup of $G$ is acylindrically arboreal.
@@
    Let $K$ be a finite normal subgroup of $G$. Then the action of $K$ on $T$ must have a fixed point $x$ by Lemma 2.9. Let $T^{\prime}$ be the convex hull of the orbit of $x$ in $T$, on which $G$ still acts non-elementarily acylindrically. Then by the normality of $K$ it must fix the entire orbit of $x$ pointwise, and hence fix all of $T^{\prime}$. The action of $G$ on $T^{\prime}$ must therefore factor through the quotient $G/K$, and thus $G/K$ is acylindrically arboreal as claimed.
@@
    This claim follows immediately from a restriction of the proof of Lemma 7.1 of [Osin13] to the case of acylindrically arboreal groups.∎
@@
[Osin13, Corollary 7.2] Let $G$ be a group acting acylindrically on a tree $T$, and let $H\cong H_{1}\times H_{2}$ be a subgroup of $G$ that decomposes as the direct product of two infinite groups. Then $H$ must act elliptically on $T$.
@@
and with an edge between any two vertices whose labels are joined by an edge of $G\cdot e$.
@@
It only remains to show that this action is non-elementary. Since $G$ cannot be virtually cyclic by assumption we need only show that the action of $G$ on $T$ has no fixed point, but if the action of $G$ on $T$ were to have a fixed vertex $v_{U}$ then the action on $T^{\prime}$ must fix the subtree $U\subseteq T^{\prime}$ setwise contradicting our assumption on $T^{\prime}$. The action of $G$ on $T$ is therefore is non-elementary acylindrical as required. ∎
@@ 3. Graph Products of Groups
Let $\Gamma=(V,E)$ be a finite simple graph, and let $\mathcal{G}=\{G_{v}\}_{v\in V}$ be a collection of groups enumerated by the vertex set of $\Gamma$. We define the graph product $\mathbb{GP}(\Gamma,\mathcal{G})$ to be the group
$\langle\mathcal{G}\mid[G_{u},G_{v}]\text{ for }\{u,v\}\in E\rangle.$
Thus $\mathbb{GP}(\Gamma,\mathcal{G})$ is the quotient of the free product of the groups $\mathcal{G}$ by the normal closure of the commutators of those pairs of groups whose labels appear in the edge set of $\Gamma$.
@@
We will use a concept of standard form for an element of a graph product $G=\mathbb{GP}(\Gamma,\mathcal{G})$ originally formulated by Green [Green1990GraphPO], although our discussion will follow that of Antolín and Minasyan [Antolinminasyan11, Section 2].
For any element $g\in G$ we can write $g$ as a word $W=(g_{1},...,g_{n})$, where $g=g_{1}\cdot\cdot\cdot g_{n}$ and each $g_{i}$ is an element of some $G_{v}\in\mathcal{G}$. We will call each $g_{i}$ a syllable of the word $W$. We say that $n$ is the length of the word $W$. We say that two syllables $g_{i}$ and $g_{j}$ of $W$ with $i<j$ can be joined together if either syllable is $1$ or if there exists some $G_{v}\in\mathcal{G}$ such that $g_{i},g_{j}\in G_{v}$ and for all $i<k<j$, $g_{k}\in G_{u_{k}}$ with $u_{k}\in N_{\Gamma}(v)$. In such a case $g_{j}$ commutes with all such $g_{k}$ in $G$, so $W$ represents the same element of $G$ as the word $(g_{1},...,g_{i-1},g_{i}g_{j},g_{i+1},...,g_{j-1},g_{j+1},...,g_{n})$, whose length is strictly smaller.
@@
[Green1990GraphPO, Theorem 3.9] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a graph product. Then every element of $G$ can be represented by a reduced word. Moreover, if two reduced words represent the same element of $G$ then one can be obtained from the other by applying a finite sequence of syllable shuffling.
Let $g\in G=\mathbb{GP}(\Gamma,\mathcal{G})$ and $W=(g_{1},...,g_{n})$ be a reduced word representing $g$. We define the length $|g|_{\Gamma}=n$ and the support of $g$ to be
$\operatorname{supp}_{\Gamma}(g)=\{v\in V(\Gamma)\mid\exists i\in\{1,...,n\}\text{ such that }g_{i}\in G_{v}\}.$
@@
Finally, we define $\operatorname{l_{V_{\Gamma}}}(g)$ and $\operatorname{f_{V_{\Gamma}}}(g)$ to be the sets of all $u\in V(\Gamma)$ such that some reduced word for $g$ ends or begins with a syllable from $G_{u}$ respectively (here f and l stand for “first” and “last”). All of these concepts are all well defined by Theorem 3.3.
@@ 3.1. Full and Parabolic Subgroups of Graph Products of Groups
This section will list without proof a collection of results of Antolin, Minasyan and Osin [Antolinminasyan11, minasyanOsin13] on the subject of full and parabolic subgroups that will provide us with the tools to study acylindrical splittings of graph products of groups.
@@
[Antolinminasyan11, Lemma 3.2] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups. Suppose that $U\subseteq V(\Gamma)$ and $g,x,y$ are some elements of $G$ such that $x$ is non-trivial, $gxg^{-1}=y$, $\operatorname{supp}_{\Gamma}(y)\subseteq U$ and $\operatorname{l_{V_{\Gamma}}}(g)\cap\operatorname{supp}_{\Gamma}(x)=\emptyset$. Then $g$ can be represented by a reduced word $(h_{1},...,h_{r},h_{r+1},...,h_{n})$, where $h_{1},...,h_{r}\in G_{U}$ and $h_{r+1},...,h_{n}\in G_{\operatorname{link}_{\Gamma}(\operatorname{supp}_{\Gamma}(x))}$.
@@
This result holds for trivial $x$ by (somewhat counter intuitively) defining the link of the empty set in a graph $\Gamma$ to be the entire vertex set. The result in such a case is completely trivial, so we have opted to exclude it so as to highlight the interesting case.
@@
[Antolinminasyan11, Lemmas 3.3, 3.4, Corollary 3.8], [minasyanOsin13, Lemma 6.4] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a graph product of groups. Then the following hold.
@@
Let $G=\mathbb{G}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups. Suppose that $P=gG_{U}g^{-1}$ with $U\subseteq V(\Gamma)$ and $g\in G$, so $P$ is a parabolic subgroup of $G$. We define the essential support $\operatorname{esupp}_{\Gamma}(P)$ to be the set $U$. This is well-defined by part (3) of Lemma 3.6.
@@ 3.2. Acylindrical Hyperbolicity of Graph Products
In this section we recall results of Minasyan and Osin which provide a condition for a graph product of groups to be acylindrically hyperbolic. Recall that a graph $\Gamma$ is irreducible if its graph theoretical complement — the graph obtained from $\Gamma$ be replacing every edge with a non-edge and every non-edge with an edge — is connected.
@@
[minasyanOsin13, Theorem 2.12] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups with $\Gamma$ irreducible. Suppose that $H\leq G$ is a subgroup that is not contained in a proper parabolic subgroup of $G$. Then $H$ is either virtually cyclic or acylindrically hyperbolic.
@@
[minasyanOsin13, Corollary 2.13] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups with $\Gamma$ irreducible. Then $G$ is either virtually cyclic or acylindrically hyperbolic.
@@
Let $T$ be the Bass–Serre associated to this splitting, and let $P$ be a path in $T$ with three edges. We may assume without loss of generality that the middle edge is labelled $G_{C}$, so there exist $g\in G_{A}-G_{C}$, $h\in G_{B}-G_{C}$ such that the other two edges in $P$ are labelled $gG_{C}$ and $hG_{C}$, as shown in Figure 1.
[…]
@@
If $gG_{C}g^{-1}\cap G_{C}$ contains only the trivial element then $\operatorname{PStab}_{G}(P)$ is trivial and we are done, so assume that there exists $y\in gG_{C}g^{-1}\cap G_{C}$ that is non-trivial. Then $y\in G_{C}$ and $y=gxg^{-1}$ for some non-trivial $x\in G_{C}$. Since $x$ is supported in $C$ we have that $\operatorname{supp}_{\Gamma}(x)\cap\operatorname{l_{V_{\Gamma}}}(g)=\emptyset$ and we can invoke Lemma 3.5 with $U=C$, which shows that $g$ is supported in $C\cup\operatorname{link}_{\Gamma}(\operatorname{supp}_{\Gamma}(x))$. Thus if $\operatorname{supp}_{\Gamma}(x)$ contains a vertex outside of $N_{\Gamma}(a)$ then $g$ is supported in $B$, contradicting that $g\in G_{A}-G_{C}$ which is disjoint from $G_{B}$. It follows that $x$ is supported in $N_{\Gamma}(a)$, and so $y\in gG_{N_{\Gamma}(a)}g^{-1}$ meaning that
@@
This last set is conjugate to $(h^{-1}gG_{N(a)}g^{-1}h)\cap G_{N(b)}$, and by Lemma 3.6 (2) there exists some $f\in G$ and $N^{\prime}\subseteq N$ such that
@@
It follows that
$|\operatorname{PStab}_{G}(P)|\leq|fG_{N^{\prime}}f^{-1}|\leq|G_{N}|,$
which is finite by assumption. Thus the action of $G$ on the Bass–Serre tree $T$ given by $G_{A}*_{G_{C}}G_{B}$ must be $(3,|G_{N}|)$-acylindrical, and so must be acylindrical by Theorem 2.17.
@@
By Example 3.9 we must then have that $\operatorname{diam}(\Gamma)<3$, so by assumption we have that that $\operatorname{diam}(\Gamma)=2$. Assume that $G$ acts acylindrically on some tree $T$, and fix the generating set $S$ of $G$ to be the union of a collection of generating sets of its vertex groups. Let $a,b\in V(\Gamma)$ be distinct vertices. We separate into two cases based on whether or not $a$ and $b$ lie in an induced $P_{3}$ subgraph of $\Gamma$.
@@
    Subcase 1, both $G_{a}$ and $G_{b}$ are finite, so $G_{\{a,b\}}$ is finite and must act elliptically.
@@
    Subcase 2, both $G_{a}$ and $G_{b}$ are infinite, so $G_{\{a,b\}}$ is the product of two infinite groups and thus acts elliptically by Lemma 2.19 as above.
@@
    Subcase 3, exactly one of $G_{a}$ and $G_{b}$ is infinite. Without loss of generality assume that $G_{a}$ is infinite and $G_{b}$ is finite. Assume for contradiction that $G_{\{a,b\}}$ contains a loxodromic element $g$. We claim that this implies that $\Gamma$ is a complete graph.
@@
    If there exists $v\in V(\Gamma)-\Lambda_{a,b}$ then $v$ must be neighbour to an element of $\operatorname{link}_{\Gamma}(\{a,b\})$ to avoid inducing a $P_{3}$ containing $a$ and $b$. However then we would have that $a,v$ is a separated pair, contradicting our initial assumption. It follows that $V(\Gamma)=\Lambda_{a,b}$, and $\Gamma$ is complete as claimed.
@@
Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups with infinite vertex groups. Then $G$ is acylindrically arboreal if and only if $\operatorname{diam}(\Gamma)\geq 3$.
@@
For the if direction the diameter being at least three implies the existence a pair of separated vertices $\{a,b\}$ by Example 3.9. The groups $G_{a}$ and $G_{b}$ are both infinite, so $G$ is not virtually cyclic and $G$ is acylindrically arboreal by Theorem 1.3.
@@
Figure 2. The underlying graph $\Gamma$ of the right angled Artin group in Proposition 1.5 (left) and its graph theoretical compliment (right).

### v6

@@ 2.3. Acylindrically Arboreal Groups
We say that a group $G$ is acylindrically arboreal if $G$ acts acylindrically non-elementarily on some tree $T$. This action will give rise to a quotient graph of groups decomposition of $G$, which we will call a non-elementary acylindrical splitting of $G$.
Any acylindrically arboreal group is acylindrically hyperbolic, as any simplicial tree is a 0-hyperbolic metric space, but working with a simplicial tree allows us to use a much more combinatorial definition.
@@
[undefao, Introduction] Let $G$ be a group acting on some tree $T$ and let $k\geq 0$ and $C>0$ be integers. We say that the action of $G$ on $T$ is $(k,C)$-acylindrical if the pointwise stabiliser of any edge path in $T$ of length at least $k$ contains at most $C$ elements.
The following theorem is essentially due to Minasyan and Osin [undefz]. We include a brief proof demonstrating how to use the cited lemma to obtain this result.
@@
[undefz, Lemma 4.2] Let $G$ be a group acting by isometries on a tree $T$. This action is acylindrical (in the sense of Definition 2.1) if and only if there exist constants $k\geq 0$ and $C\geq 1$ such that the action of $G$ on $T$ is $(k,C)$-acylindrical.
@@
For the other direction assume that the action of $G$ on $T$ is $(k,C)$-acylindrical for some constants $k$ and $C$. Let $\epsilon>0$, and set
$R(\epsilon)=k+2\epsilon+6\text{, }N(\epsilon)=2(2\epsilon+1)C.$
@@
so it follows that $d_{T}(u,v)\geq k$ and $d_{T}(\{u,v\},\{x,y\})>\epsilon,$ and we can invoke [undefz, Lemma 4.2] to see that the pointwise quasi-stabiliser $\operatorname{PQStab}_{G}^{\epsilon}(\{x,y\})$ is contained in at most $2(2\epsilon+1)$ cosets of $\operatorname{PStab}_{G}(\{u,v\})$, the pointwise stabiliser of $\{u,v\}$. By construction the geodesic path between $u$ and $v$ must use at least $k$ edges and so $|\operatorname{PStab}_{G}(\{u,v\})|\leq C$ by definition of definition of $k$ and $C$. It follows that $|\operatorname{PQStab}_{G}^{\epsilon}(\{x,y\})|\leq 2(2\epsilon+1)C=N(\epsilon)$. Thus our action was acylindrical with constants $R(\epsilon)$ and $N(\epsilon)$. ∎
The class of acylindrically arboreal groups has the following set of inheritance properties, claim (3) of which follows immediately from a restriction of [undefaf, Lemma 7.1] to actions on trees.
@@
    [undefaf, Lemma 7.1] Let $H\leq G$ be s-normal, i.e. for all $g\in G$, $gHg^{-1}\cap H$ is infinite. Then $H$ is acylindrically arboreal. In particular, any finite index subgroup of $G$ is acylindrically arboreal.
@@
    Let $K$ be a finite normal subgroup of $G$. Then the action of $K$ on $T$ must have a fixed point $x$ by Lemma 2.9. Let $T^{\prime}$ be the convex hull of the orbit of $x$ in $T$ under the action of $G$, on which $G$ still acts non-elementarily acylindrically. Then by the normality of $K$ it must fix the entire orbit of $x$ pointwise, and hence fix all of $T^{\prime}$. The action of $G$ on $T^{\prime}$ must therefore factor through the quotient $G/K$, and thus $G/K$ is acylindrically arboreal as claimed.
@@
    This claim follows immediately from a restriction of the proof of [undefaf, Lemma 7.1] to the case of acylindrically arboreal groups.∎
@@
[undefaf, Corollary 7.2] Let $G$ be a group acting acylindrically on a tree $T$, and let $H\cong H_{1}\times H_{2}$ be a subgroup of $G$ that decomposes as the direct product of two infinite groups. Then $H$ must act elliptically on $T$.
@@
and an edge between any two vertices whose labels are joined by an edge of $G\cdot e$.
@@
It only remains to show that this action is non-elementary. Since $G$ cannot be virtually cyclic by assumption we need only show that the action of $G$ on $T$ has no fixed point, but if the action of $G$ on $T$ were to have a fixed vertex $v_{U}$ then the action on $T^{\prime}$ must fix the subtree $U\subseteq T^{\prime}$ setwise contradicting our assumption on $T^{\prime}$. The action of $G$ on $T$ is therefore non-elementary acylindrical as required. ∎
@@ 3. Graph Products of Groups
Let $\Gamma=(V,E)$ be a finite simple graph, and let $\mathcal{G}=\{G_{v}\}_{v\in V}$ be a collection of groups enumerated by the vertex set of $\Gamma$. We define the graph product $\mathbb{GP}(\Gamma,\mathcal{G})$ to be the group $\langle\mathcal{G}\mid[G_{u},G_{v}]$ for $\{u,v\}\in E\rangle$. Thus $\mathbb{GP}(\Gamma,\mathcal{G})$ is the quotient of the free product of the groups $\mathcal{G}$ by the normal closure of the commutators of those pairs of groups whose labels appear in the edge set of $\Gamma$.
@@
We will use a concept of standard form for an element of a graph product $G=\mathbb{GP}(\Gamma,\mathcal{G})$ originally formulated by Green [undefm], although our discussion will follow that of Antolín and Minasyan [undef, Section 2].
For any element $g\in G$ we can write $g$ as a word $W=(g_{1},...,g_{n})$, where $g=g_{1}\cdot\cdot\cdot g_{n}$ and each $g_{i}$ is an element of some $G_{v}\in\mathcal{G}$. We will call each $g_{i}$ a syllable of the word $W$. We say that $n$ is the length of the word $W$. We say that two syllables $g_{i}$ and $g_{j}$ of $W$ with $i<j$ can be joined together if either syllable is $1$ or if there exists some $G_{v}\in\mathcal{G}$ such that $g_{i},g_{j}\in G_{v}$ and for all $i<k<j$, $g_{k}\in G_{u_{k}}$ with $u_{k}\in N_{\Gamma}(v)$. In such a case $g_{j}$ commutes with all such $g_{k}$ in $G$, so $W$ represents the same element of $G$ as the word
$(g_{1},...,g_{i-1},g_{i}g_{j},g_{i+1},...,g_{j-1},g_{j+1},...,g_{n})$
, whose length is strictly smaller.
@@
[undefm, Theorem 3.9] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a graph product. Then every element of $G$ can be represented by a reduced word. Moreover, if two reduced words represent the same element of $G$ then one can be obtained from the other by applying a finite sequence of syllable shuffling.
Let $g\in G=\mathbb{GP}(\Gamma,\mathcal{G})$ and let $W=(g_{1},...,g_{n})$ be a reduced word representing $g$. We define the length $|g|_{\Gamma}=n$ and the support of $g$ to be
$\operatorname{supp}_{\Gamma}(g)=\{v\in V(\Gamma)\mid\text{there exists }i\in\{1,...,n\}\text{ such that }g_{i}\in G_{v}\}.$
@@
Finally, we define $\operatorname{l_{V_{\Gamma}}}(g)$ and $\operatorname{f_{V_{\Gamma}}}(g)$ to be the sets of all $u\in V(\Gamma)$ such that some reduced word for $g$ ends or begins with a syllable from $G_{u}$ respectively (here f and l stand for “first” and “last”). All of these concepts are well defined by Theorem 3.3.
@@ 3.1. Full and Parabolic Subgroups of Graph Products of Groups
This section will list without proof a collection of results of Antolín, Minasyan and Osin [undef, undefz] on the subject of full and parabolic subgroups that will provide us with the tools to study acylindrical splittings of graph products of groups.
@@
[undef, Lemma 3.2] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups. Suppose $U\subseteq V(\Gamma)$ and $g,x,y$ are some elements of $G$ such that $x$ is non-trivial, $gxg^{-1}=y$, $\operatorname{supp}_{\Gamma}(y)\subseteq U$ and $\operatorname{l_{V_{\Gamma}}}(g)\cap\operatorname{supp}_{\Gamma}(x)=\emptyset$. Then $g$ can be represented by a reduced word $(h_{1},...,h_{r},h_{r+1},...,h_{n})$, where $h_{1},...,h_{r}\in G_{U}$ and $h_{r+1},...,h_{n}\in G_{\operatorname{link}_{\Gamma}(\operatorname{supp}_{\Gamma}(x))}$.
@@
This result holds for trivial $x$ by (somewhat counter-intuitively) defining the link of the empty set in a graph $\Gamma$ to be the entire vertex set. The result in such a case is completely trivial, so we have opted to exclude it so as to highlight the interesting case.
@@
[undef, Lemmas 3.3, 3.4, Corollary 3.8], [undefz, Lemma 6.4] Suppose that $G=\mathbb{GP}(\Gamma,\mathcal{G})$ is a graph product of groups. Then the following hold.
@@
Let $G=\mathbb{G}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups. Suppose that $P=gG_{U}g^{-1}$ with $U\subseteq V(\Gamma)$ and $g\in G$, so $P$ is a parabolic subgroup of $G$. We define the essential support $\operatorname{esupp}_{\Gamma}(P)$ to be the set $U$. This is well-defined Lemma 3.6(3).
@@ 3.2. Acylindrical Hyperbolicity of Graph Products
In this section we recall results of Minasyan and Osin which provide a condition for a graph product of groups to be acylindrically hyperbolic. Recall that a graph $\Gamma$ is irreducible if its graph theoretical complement — the graph obtained from $\Gamma$ by replacing every edge with a non-edge and every non-edge with an edge — is connected.
@@
[undefz, Theorem 2.12] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups with $\Gamma$ irreducible. Suppose that $H\leq G$ is a subgroup that is not contained in a proper parabolic subgroup of $G$. Then $H$ is either virtually cyclic or acylindrically hyperbolic.
@@
[undefz, Corollary 2.13] Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups with $\Gamma$ irreducible. Then $G$ is either virtually cyclic or acylindrically hyperbolic.
@@
Let $T$ be the Bass–Serre tree associated to this splitting, and let $P$ be a path in $T$ with three edges. We may assume without loss of generality that the middle edge is labelled $G_{C}$, so there exist $g\in G_{A}-G_{C}$, $h\in G_{B}-G_{C}$ such that the other two edges in $P$ are labelled $gG_{C}$ and $hG_{C}$, as shown in Figure 1.
$G_{A}$$G_{B}$$gG_{B}$$hG_{A}$$G_{C}$$gG_{C}$$hG_{C}$
@@
If $gG_{C}g^{-1}\cap G_{C}$ contains only the trivial element then $\operatorname{PStab}_{G}(P)$ is trivial and we are done, so assume that there exists $y\in gG_{C}g^{-1}\cap G_{C}$ that is non-trivial. Then $y\in G_{C}$ and $y=gxg^{-1}$ for some non-trivial $x\in G_{C}$. Since $x$ is supported in $C$ we have that $\operatorname{supp}_{\Gamma}(x)\cap\operatorname{l_{V_{\Gamma}}}(g)=\emptyset$ and we can invoke Lemma 3.5 with $U=C$, which shows that $g$ is supported in $C\cup\operatorname{link}_{\Gamma}(\operatorname{supp}_{\Gamma}(x))$. Thus if $\operatorname{supp}_{\Gamma}(x)$ contains a vertex outside of $N_{\Gamma}(a)$ then $g$ is supported in $B$, contradicting the fact that $g\in G_{A}-G_{C}$ which is disjoint from $G_{B}$. It follows that $x$ is supported in $N_{\Gamma}(a)$, and so $y\in gG_{N_{\Gamma}(a)}g^{-1}$ meaning that
@@
This last set is conjugate to $(h^{-1}gG_{N(a)}g^{-1}h)\cap G_{N(b)}$, and by Lemma 3.6 (2) there exist some $f\in G$ and $N^{\prime}\subseteq N$ such that
@@
It follows that $|\operatorname{PStab}_{G}(P)|\leq|fG_{N^{\prime}}f^{-1}|\leq|G_{N}|,$ which is finite by assumption. Thus the action of $G$ on the Bass–Serre tree $T$ given by $G_{A}*_{G_{C}}G_{B}$ must be $(3,|G_{N}|)$-acylindrical, and so must be acylindrical by Theorem 2.17.
@@
By Example 3.9 we must then have that $\operatorname{diam}(\Gamma)<3$, so by assumption we have that $\operatorname{diam}(\Gamma)=2$. Assume that $G$ acts acylindrically on some tree $T$, and fix the generating set $S$ of $G$ to be the union of a collection of generating sets of its vertex groups. Let $a,b\in V(\Gamma)$ be distinct vertices. We separate into two cases based on whether or not $a$ and $b$ lie in an induced $P_{3}$ subgraph of $\Gamma$.
@@
    Both $G_{a}$ and $G_{b}$ are finite, so $G_{\{a,b\}}$ is finite and must act elliptically.
@@
    Both $G_{a}$ and $G_{b}$ are infinite, so $G_{\{a,b\}}$ is the product of two infinite groups and thus acts elliptically by Lemma 2.19 as above.
@@
    Exactly one of $G_{a}$ and $G_{b}$ is infinite. Without loss of generality assume that $G_{a}$ is infinite and $G_{b}$ is finite. Assume for contradiction that $G_{\{a,b\}}$ contains a loxodromic element $g$. We claim that this implies that $\Gamma$ is a complete graph.
@@
    Assume for contradiction that there exists $v\in V(\Gamma)-\Lambda_{a,b}$. We may assume without loss of generality that $v$ is neighbour to some element of $\Lambda_{a,b}$ by the fact that $\operatorname{diam}(\Gamma)=2$. If $v$ is a neighbour of one of $a$ or $b$, then it must also be a neighbour of the other to avoid inducing a $P_{3}$ on $v,a,b$, and so $v\in\operatorname{link}_{\Gamma}(\{a,b\})$ contradicting our choice of $v$. It follows that $v$ is a neighbour of some element of $\operatorname{link}_{\Gamma}(\{a,b\})$. However, we then have that $a,v$ is a separated pair as $\operatorname{link}_{\Gamma}(\{a,v\})\subseteq\operatorname{link}_{\Gamma}(a)\subseteq\operatorname{link}_{\Gamma}(\{a,b\})$ which corresponds to a finite full subgroup, providing the desired contradiction. It follows that $V(\Gamma)=\Lambda_{a,b}$, and $\Gamma$ is complete as claimed.
@@
Let $G=\mathbb{GP}(\Gamma,\mathcal{G})$ be a non-degenerate graph product of groups with infinite vertex groups. Then we have that $G$ is acylindrically arboreal if and only if $\operatorname{diam}(\Gamma)\geq 3$.
@@
For the if direction the diameter being at least three implies the existence of a pair of separated vertices $\{a,b\}$ by Example 3.9. The groups $G_{a}$ and $G_{b}$ are both infinite, so $G$ is not virtually cyclic and $G$ is acylindrically arboreal by Theorem 1.3.
@@
Figure 2. The underlying graph $\Gamma$ of the right angled Artin group in Proposition 1.5 (left) and its graph theoretical complement (right).